OpenStudy (anonymous):

Which property is shown? Additive Identity Property Multiplicative Inverse Property Associate Property of Addition Additive Inverse Property A. Maybe?

OpenStudy (anonymous):

Solve: 4/5+(-4/5)=0

OpenStudy (anonymous):

I'm thinking it might be A but not sure.

OpenStudy (anonymous):

I think its d because the additive identity would just be 4/5+0=0 which is wrong and theres no multiplication so it cant be b either

OpenStudy (anonymous):

Is that the final answer, d?

OpenStudy (anonymous):

@WiredCardinal08

OpenStudy (anonymous):

Thats what I would put

OpenStudy (anonymous):

Dont hate me if I'm wrong tho lol

OpenStudy (anonymous):

Dang, I'm nervous man lol. How about this? Which property is shown? Multiplicative Identity Distributive Property Multiplicative Inverse Property Commutative Property of Multiplication

jimthompson5910 (jim_thompson5910):

The additive property is the idea that adding 0 to any number gives you that same exact number ex: 7+0 = 7

jimthompson5910 (jim_thompson5910):

The additive inverse property is a related idea that says if you add a number with its opposite, then you get 0 ex: 7 + (-7) = 0

jimthompson5910 (jim_thompson5910):

additive identity property*

OpenStudy (anonymous):

Alright, how about the second question?

jimthompson5910 (jim_thompson5910):

you listed the choices, but forgot the question

OpenStudy (anonymous):

Sorry about that. 2(x-4)=2*x-2*4

jimthompson5910 (jim_thompson5910):

what are your thoughts on it?

OpenStudy (anonymous):

Give me a minute on this one.

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

Distributive Property

jimthompson5910 (jim_thompson5910):

very good

OpenStudy (anonymous):

Alright guys thanks for helping out.